### v2

# GS-KAN: Parameter-Efficient Kolmogorov-Arnold Networks via Sprecher-Type Shared Basis Functions

Abstract
The Kolmogorov-Arnold representation theorem offers a theoretical alternative to Multi-Layer Perceptrons (MLPs) by placing learnable univariate functions on edges rather than nodes. While recent implementations such as Kolmogorov-Arnold Networks (KANs) demonstrate high approximation capabilities, they suffer from significant parameter inefficiency due to their reliance on unique parameterizations for every network edge. In this work, we propose GS-KAN (Generalized Sprecher-KAN), a lightweight architecture inspired by David Sprecher’s refinement of the superposition theorem. GS-KAN constructs unique edge functions by applying learnable linear transformations to a single learnable, shared parent function per layer. We evaluate GS-KAN against existing KAN architectures and MLPs across synthetic function approximation, tabular data regression, and image classification tasks. Our results demonstrate that GS-KAN outperforms both MLPs and standard KAN baselines on continuous function approximation while maintaining superior parameter efficiency. Additionally, GS-KAN achieves competitive performance with existing KAN architectures on tabular regression and outperforms MLPs on high-dimensional classification tasks. Crucially, the proposed architecture enables the deployment of KAN-based architectures in high-dimensional regimes under strict parameter constraints, a setting where standard implementations are typically infeasible due to parameter explosion. The source code is available at https://github.com/rambamn48/gs-impl.
I Introduction
The Multi-Layer Perceptron (MLP) has long served as the foundational building block of deep learning. By stacking linear transformations followed by fixed element-wise non-linearities (e.g., ReLU), MLPs leverage the Universal Approximation Theorem [1, 2] to model complex data manifolds. However, the recent introduction of Kolmogorov-Arnold Networks (KANs) [3] challenges this paradigm by placing learnable univariate functions on the edges of the network. KANs have shown promise in interpretability and data efficiency, particularly for low-dimensional scientific discovery tasks.
However, standard KANs face a critical scalability bottleneck. The reliance on unique function parameterization for every edge leads to a parameter complexity of $O(C\cdot N_{in}\cdot N_{out})$, where $C$ is the number of parameters for the learned functions. For high-dimensional inputs, such as flattened images, this results in an explosion of parameters that renders standard KANs impractical or prohibitive to deploy under strict memory budgets. While recent variants like Wav-KAN [4] attempt to address this by substituting splines with fixed wavelets, they often sacrifice the learnable adaptability of the basis function.
In this paper, we revisit the theoretical foundations laid by Sprecher [5] to address this inefficiency. While the original KAN architecture [3] leverages the Kolmogorov-Arnold representation theorem [6], it does not fully exploit Sprecher’s subsequent refinement, which demonstrated that the inner functions can be constructed as transformed versions of a single univariate function. Historically, direct implementations of such “Sprecher networks” were hindered by the pathological, fractal nature of the exact functions required by the theorem, rendering them unsuitable for gradient descent (see, e.g., [7]).
We propose GS-KAN (Generalized Sprecher-KAN), which relaxes these strict conditions. Instead of attempting to model exact fractal functions, we employ smooth, learnable B-splines combined with generalized coefficients. Unlike standard KANs which maintain unique splines for every edge, GS-KAN maintains one learnable master function per layer. This reduces the parameter complexity, effectively aligning it with MLP efficiency while retaining the powerful inductive bias of splines.
Our contributions are as follows:
  • Parameter Efficiency: We propose a memory-efficient formulation using layer-wise shared learnable B-splines, adapted via learnable linear parameters. This architecture significantly reduces the parameter footprint compared to standard KANs.
  • Scalability: By decoupling the basis definition from the network edges, GS-KAN enables the use of spline-based architectures on high-dimensional inputs (e.g., image vectors), overcoming the parameter bottlenecks of previous implementations.
  • Empirical Validation: We demonstrate that GS-KAN is highly versatile. It outperforms MLPs and KAN architectures on high-frequency function approximation and matches or exceeds existing KAN variants on tabular benchmarks. Furthermore, we show that GS-KAN can successfully classify high-dimensional image vectors (Fashion-MNIST) with fewer parameters than MLPs.
A recent concurrent study by [8] also explores the application of Sprecher’s theorem for parameter efficiency. Their approach focuses on a strict realization of Sprecher’s specific summation formula using shared weight vectors. In contrast our work builds upon a variation of the theorem and from this basis, we propose a generalized architecture where we relax the fixed constant constraints to construct unique edge functions via learnable per-edge linear transformations. This distinction allows our method to retain the memory benefits of the theorem while adopting the flexible connectivity typical of modern deep learning layers.
This work extends the initial investigations documented in the author’s Bachelor’s thesis [9]. Although the core concept of leveraging Sprecher’s theorem originated there, the present work introduces a novel, fully generalized architecture explicitly designed to support arbitrary input-output mappings—including multi-output regression and classification tasks—accompanied by improved initialization strategies and deeper empirical analysis.
[…]
Figure 1: Edge-Node Computation Comparison. Left: Standard KAN learns a unique function $\phi_{q,p}$ for every connection. Right: GS-KAN learns a single shared basis $\psi$ and adapts it to each edge via learnable scalars $\lambda$ and shifts $\epsilon$. The node $y_{q}$ aggregates the results.
II Theoretical Background
This section outlines the mathematical framework underpinning neural networks. By composing elementary operations into layers and stacking them sequentially, we establish the deep architecture required for complex modeling. The following theorems describe the capabilities of this hierarchical structure.
II-A Universal Approximation Theorem (MLP)
The Multi-Layer Perceptron is grounded in the Universal Approximation Theorem, initially formulated by [1] and [2]. This theorem guarantees that MLPs can act as universal function approximators under general conditions.
Theorem 1 (Universal Approximation Theorem).
Let $\sigma(\cdot)$ be a fixed, non-linear activation function. Any continuous function $f:\mathbb{R}^{n}\to\mathbb{R}$ can be approximated to arbitrary accuracy by a finite linear combination of the form:
$f(\mathbf{x})\approx\sum_{i=1}^{N}v_{i}\sigma\left(\sum_{j=1}^{n}w_{ij}x_{j}+b_{i}\right)$ (1)
where $N$ is the number of hidden neurons, and $w_{ij},v_{i},b_{i}$ are learnable parameters.
In this formulation, the non-linearity $\sigma$ (e.g., ReLU or Sigmoid) is fixed and applied after the linear aggregation of inputs. This stands in contrast to KANs, where the non-linearities are learnable univariate functions on the edges.
II-B Kolmogorov-Arnold Representation (KAN)
In contrast to the fixed nodes of MLPs, Kolmogorov [6] proved that any continuous multivariate function on a compact domain $0,1]^{n}$ can be represented exactly as a superposition of continuous univariate functions.
Theorem 2 (Kolmogorov-Arnold Representation).
For any continuous function $f:[0,1]^{n}\to\mathbb{R}$, there exist continuous univariate functions $\Phi_{q}$ and $\phi_{q,p}$ such that:
$f(x_{1},\dots,x_{n})=\sum_{q=0}^{2n}\Phi_{q}\left(\sum_{p=1}^{n}\phi_{q,p}(x_{p})\right)$ (2)
The recently proposed KAN architecture [3] parameterizes these univariate functions $\phi_{q,p}$ as B-splines. While this allows for highly accurate approximation, it requires learning a unique spline (and its associated grid points) for every edge $(p,q)$ in the network.
II-C Sprecher’s Refinement
Sprecher [5] significantly refined the representation theory by demonstrating that distinct functions for each branch are not required. Instead, the theorem guarantees that a multivariate function can be represented using single continuous functions $\chi$ and $\psi$ through a specific construction involving scaling and translation.
Theorem 3 (Sprecher, 1965).
For any continuous function $f:[0,1]^{n}\to\mathbb{R}$, there exist real constants $\lambda,\epsilon$ and a continuous function $\chi:\mathbb{R}\to\mathbb{R}$ such that:
$f(\mathbf{x})=\sum_{q=0}^{2n}\chi\left(\sum_{p=1}^{n}\lambda^{p\cdot q}\psi(x_{p}+\epsilon q)\right)$ (3)
Here, $\psi$ is a universal function shared across all terms. Crucially, the outer function $\chi$ is also shared (independent of $q$), acting on the aggregated summation.
II-D Related Efficient Architectures
Recent works have attempted to streamline KANs. Notably, Wav-KAN [4] replaces B-splines with wavelets, defined as $\psi_{w,s,t}(x)=w\psi(\frac{x-t}{s})$. While Wav-KAN shares the philosophy of transforming a basis function, it typically relies on a fixed analytical mother wavelet (e.g., Mexican Hat). Our approach differs by making the basis function $\psi$ itself a learnable B-spline, optimizing both the transformation and the shape of the function simultaneously.
III Methodology
To translate Sprecher’s theoretical framework into a flexible deep learning network, we introduce key relaxations to the strict formulation. While the theorem postulates the existence of specific fixed constants to ensure exact representation, enforcing such rigid values restricts the model’s adaptability during gradient-based optimization.
III-A Generalized Representation
Our architecture, GS-KAN, relaxes the theorem’s fixed structure into flexible, learnable components. Specifically, we introduce two primary generalizations:
  1. 1.
    Per-Edge Weighting ($\lambda^{p\cdot q}\to\lambda_{p,q}$): Instead of the fixed power term $\lambda^{p\cdot q}$ derived from the theorem, we assign a unique, learnable weight $\lambda_{p,q}$ to every edge connecting input dimension $p$ to hidden node $q$.
  2. 2.
    Independent Translation ($\epsilon\cdot q\to\epsilon_{q}$): We replace the fixed shift term (typically scaled by index $q$) with independent learnable parameters $\epsilon_{q}$, allowing for free shifts of the layer functions.
This flexibility allows the network to dynamically discover the optimal separation and alignment of terms in the high-dimensional feature space. Rather than being confined to a predetermined structure, GS-KAN leverages these decoupled parameters to better capture the underlying data distribution.
Formally, we define the mapping between an input layer of width $N_{in}$ and an output layer of width $N_{out}$ as follows. Let $p\in\{1,\dots,N_{in}\}$ denote the index of the sending node, and $q\in\{1,\dots,N_{out}\}$ denote the receiving node. The value of node $y_{q}$ is given by:
$y_{q}=\sum_{p=1}^{N_{in}}\lambda_{p,q}\cdot\psi_{l}\left(x_{p}+\epsilon_{q}\right)$ (4)
where $\lambda_{p,q}$ represents a learnable edge weight and $\epsilon_{q}$ is a learnable node-specific bias (translation). See Figure 1 for an illustration of this construction.
Distinction from MLPs.
It is crucial to distinguish this topology from the standard Multi-Layer Perceptron (MLP). MLPs rely on the Universal Approximation Theorem by applying a fixed non-linearity $\sigma$ after a linear summation: $y_{q}=\sigma(\sum w_{qp}x_{p}+b_{q})$. In contrast, GS-KAN adheres to the Kolmogorov-Arnold topology where learnable non-linearities are applied on the edges (pre-summation). While Standard KANs use unique functions $\phi_{qp}$ for every edge, GS-KAN constrains these to be linear transformations of a single shared basis $\psi_{l}$. This maintains the KAN topology while drastically reducing parameter count.
III-B Network Architecture & Implementation
We propose a deep architecture constructed by stacking the generalized Sprecher layers described above.
Shared Learnable B-Splines.
For each layer $l$, we parameterize the shared basis function $\psi_{l}(t)$ using (cubic) B-splines. Unlike wavelets which have a fixed shape (e.g., Mexican Hat), B-splines are fully learnable, allowing the network to discover optimal activation shapes from data. The spline is defined by a set of learnable coefficients $c$ and a knot vector $K$.
Fixed-Domain Strategy.
A known challenge in KAN training is the need for adaptive grids when feature distributions drift outside the spline’s definition range. To eliminate the computational overhead of grid re-calculation, we employ a static domain strategy:
  1. 1.
    Static Grid Definition: We fix the definition domain of the B-splines to a static interval [G,G][-G,G].
  2. 2.
    Input-Only Normalization: We provide a method for standard normalization only to the raw input data, if needed, to ensure it falls within the active domain at initialization.
  3. 3.
    Implicit Adaptation via Batch Optimization: For latent layers, we rely on the learnable parameters to naturally map feature distributions into the spline’s active domain. While individual samples may occasionally fall outside the grid (yielding zero local gradients), the use of batch-based optimization ensures that the aggregate gradient remains informative. This allows the network to dynamically adjust the scaling and shifting parameters to keep the majority of activations within the valid range during training.
This approach simplifies the training loop while maintaining robustness against outliers.
III-C Parameter Complexity Analysis
The primary motivation for GS-KAN is parameter efficiency. Consider a layer with $N_{in}$ inputs and $N_{out}$ outputs. Let $C$ be the number of parameters per learned function (e.g., B-splines coefficients). For simplicity, we omit linear bias/shift terms ($N_{out}$) for both MLP and GS-KAN in this comparison.
  • Standard KAN: Requires a unique function for every edge. Parameter count $\approx N_{in}\times N_{out}\times C$.
  • MLP: Requires one weight per edge. Parameter count $\approx N_{in}\times N_{out}$.
  • GS-KAN: Requires one weight matrix ($\lambda$) and one shared spline ($\psi$). Parameter count $\approx N_{in}\times N_{out}+C$.
Since $C$ typically is negligible compared to the weight matrix ($N_{in}\times N_{out}$), GS-KAN achieves the same asymptotic parameter complexity as an MLP ($O(N_{in}\cdot N_{out})$), while retaining the mathematical properties of Kolmogorov-Arnold networks. This contrasts with Standard KANs, which are fundamentally heavier ($O(C\cdot N_{in}\cdot N_{out})$).
IV Experiments
We evaluate the performance and parameter efficiency of GS-KAN across three distinct domains: synthetic function approximation, real-world tabular regression, and high-dimensional image classification.
Baselines & Implementation.
We compare GS-KAN against three primary baselines to assess its relative efficiency:
  • MLP: A standard fully connected Multi-Layer Perceptron. We use SiLU activations for function approximation tasks (to favor smoothness) and ReLU for tabular/image tasks.
  • Std-KAN: The standard Kolmogorov-Arnold Network, evaluated using the optimized efficient-kan library.
  • Wav-KAN: A wavelet-based KAN implementation using the Mexican Hat wavelet.
Training Protocol.
All models are implemented in PyTorch and trained using the Adam optimizer. Cubic B-Splines were utilized in the GS-KAN models. To ensure statistical robustness, every experiment is repeated across three distinct initialization seeds. We report the best test metric (MSE or Accuracy) achieved during training to capture the peak potential of each architecture, simulating an optimal early-stopping strategy.
IV-A Synthetic Function Approximation
To evaluate the capability of GS-KAN to model complex, high-frequency non-linear interactions under strict parameter constraints, we conducted a regression benchmark using a 2D crossed-wave function. Here a detailed description of the experimental setup is provided.
Setup.
The target function is defined as:
$f(x,y)=\sin(3\pi x)\cdot\cos(3\pi y),\quad(x,y)\in[-1,1]^{2}$ (5)
To simulate realistic conditions, additive Gaussian noise $\epsilon\sim\mathcal{N}(0,0.01^{2})$ was introduced to the targets. The theoretical lower bound for the Mean Squared Error (MSE) is thus $1.0\times 10^{-4}$. The dataset consists of 4,096 samples trained for 150 epochs. The knot interval for the GS-KAN B-splines was set to [-1,1].
Results.
The comparison was conducted in two parameter regimes: Nano ($\approx 200$ params) and Micro ($\approx 500$ params). The specific architectures (depth and width) were adjusted to match the parameter counts across models. Results are presented in Table I and an illustration of the GS-KAN (micro) prediction is found in Figure 2.
Figure 2: Function Approximation Visualization. Left: The ground truth function $f(x,y)=\sin(3\pi x)\cdot\cos(3\pi y)$ (noiseless). Right: The approximation generated by GS-KAN (Micro configuration, $\approx 500$ parameters). Notably, the model effectively captures high-frequency structures and sharp peaks despite strict parameter constraints, a regime where MLPs typically suffer from over-smoothing.
TABLE I: Function Approximation Results. Comparison of the best test MSE (Mean $\pm$ Std). The Architecture column denotes $N_{in},N_{hidden},\dots,N_{out}]$. Resolution refers to Knots ($K$) for GS-KAN and Grid Size ($G$) for Std-KAN.
Model Architecture Res. Params Best Test MSE
(Mean $\pm$ Std)
Nano Regime ($\approx 200$ Params)
MLP $[2,12,12,1]$ - 205 $2.55\times 10^{-1}\pm 9.08\times 10^{-3}$
Std-KAN $[2,5,1]$ $G=8$ 195 $1.17\times 10^{-3}\pm 3.07\times 10^{-4}$
Wav-KAN $[2,7,7,1]$ - 210 $1.57\times 10^{-2}\pm 3.84\times 10^{-3}$
GS-KAN $\mathbf{[2,10,10,1]}$ $\mathbf{K=20}$ 190 $\mathbf{8.40\times 10^{-4}\pm 7.56\times 10^{-4}}$
Micro Regime ($\approx 500$ Params)
MLP $[2,20,20,1]$ - 501 $2.55\times 10^{-1}\pm 9.35\times 10^{-3}$
Std-KAN $[2,5,5,1]$ $G=7$ 480 $4.41\times 10^{-4}\pm 1.28\times 10^{-5}$
Wav-KAN $[2,12,11,1]$ - 501 $3.01\times 10^{-3}\pm 6.92\times 10^{-4}$
GS-KAN $\mathbf{[2,16,16,1]}$ $\mathbf{K=50}$ 478 $\mathbf{1.65\times 10^{-4}\pm 3.58\times 10^{-5}}$
00footnotetext: For GS-KAN, the number of coefficients for each function is $N_{c}=K-d-1$, whereas for Standard KAN, it is $N_{c}=G+d$, where $d=3$ is the spline degree.
Analysis.
GS-KAN outperforms all baselines in both regimes, despite utilizing the fewest learnable parameters. Notably, the MLP baseline fails to capture the target structure ($MSE\approx 0.25$) even with SiLU activations, highlighting the inherent difficulty networks face in modelling high-frequency periodicities. In the Micro regime, GS-KAN approaches the theoretical noise floor ($1.65\times 10^{-4}$), outperforming Standard KAN by a factor of $\approx 2.6$. This efficiency is driven by the shared-basis architecture, which decouples grid resolution from model size. Consequently, GS-KAN can employ significantly deeper and wider structures and higher knot counts (e.g., K=50K=50)—while maintaining the parameter budget.
IV-B Tabular Regression (California Housing)
We evaluated the models on the California Housing dataset to assess performance on noisy, real-world tabular data ($D_{in}=8$).
Setup.
The dataset contains 20,640 samples. Inputs were standardized using z-score normalization. All models were trained for 100 epochs. The knot domain was set to [3,3][-3,3].
Results.
We benchmarked performance across three parameter regimes. The specific architectures (depth and width) were adjusted to match the parameter counts across models. Results are shown in Table II.
TABLE II: California Housing Results. Comparison of Best Test MSE (Mean $\pm$ Std). The best result in each regime is highlighted in bold. Note that GS-KAN outperforms the MLP baseline in all regimes.
Model Architecture Res. Params Best Test MSE
(Mean $\pm$ Std)
Nano Regime ($\approx 200$ Params)
MLP $[8,10,10,1]$ - 211 $0.2939\pm 0.0056$
Std-KAN $[8,2,1]$ $G=7$ 216 $0.2934\pm 0.0002$
Wav-KAN $[8,5,5,1]$ - 210 $\mathbf{0.2847\pm 0.0057}$
GS-KAN $[8,8,8,1]$ $K=20$ 204 $0.2901\pm 0.0063$
Micro Regime ($\approx 600$ Params)
MLP $[8,20,20,1]$ - 621 $0.2734\pm 0.0044$
Std-KAN $[8,5,1]$ $G=8$ 585 $0.2521\pm 0.0036$
Wav-KAN $[8,11,10,1]$ - 624 $\mathbf{0.2512\pm 0.0077}$
GS-KAN $[8,18,18,1]$ $K=20$ 574 $0.2587\pm 0.0052$
Small Regime ($\approx 2000$ Params)
MLP $[8,40,40,1]$ - 2041 $0.2691\pm 0.0047$
Std-KAN $[8,8,8,1]$ $G=10$ 2040 $\mathbf{0.2294\pm 0.0032}$
Wav-KAN $[8,22,21,1]$ - 1977 $0.2436\pm 0.0003$
GS-KAN $[8,38,38,1]$ $K=40$ 1974 $0.2358\pm 0.0082$
Analysis.
GS-KAN demonstrates robust scaling properties, consistently outperforming the MLP baseline across all tested parameter regimes. While the MLP exhibits saturation, GS-KAN effectively captures high-frequency interactions. Moreover, GS-KAN remains highly competitive with state-of-the-art variants like Std-KAN and Wav-KAN, frequently matching or exceeding their performance. For instance, in the parameter-constrained Small regime GS-KAN achieves an MSE of $0.236$, surpassing Wav-KAN ($0.244$) and outperforms Std-KAN in the Nano regime.
IV-C High-Dimensional Classification (Fashion-MNIST)
To assess the versatility of GS-KAN beyond regression tasks, we evaluated its performance on a high-dimensional classification problem using the Fashion-MNIST dataset. This task tests the model’s ability to handle large input vectors ($D_{in}=784$) and extract semantic features without the inductive bias of convolutional layers.
Setup.
The dataset consists of 60,000 training images and 10,000 test images across 10 classes. Inputs were flattened to vectors of size 784 and normalized to [1,1][-1,1]. Both models were trained for a fixed budget of 20 epochs using Cross Entropy Loss. We employed a parameter constraint of $\approx 12.5k$ parameters, forcing the models to learn efficient representations with minimal capacity. We utilized B-splines with a resolution of $K=60$ knots and a domain of [3,3][-3,3]. Although the input pixels are normalized to [1,1][-1,1], the learnable parameters in the GS-KAN layers may scale or shift activations beyond the unit interval. Extending the grid ensures that these transformed features remain within the active region of the spline basis, preventing gradient saturation.
Results.
We compare GS-KAN against a standard MLP with ReLU activations. To ensure a rigorous comparison, the MLP was configured with a slightly larger parameter budget ($5\%$) than the GS-KAN model.
TABLE III: Classification Results (Fashion-MNIST). Comparison of Best Test Accuracy with a restricted parameter budget ($\approx 12.5k$). The MLP baseline was granted a larger parameter count to test the efficiency of GS-KAN.
Model Architecture Params Best Test Acc
(Mean $\pm$ Std)
MLP $[784,15,20,10]$ 13,002 $86.00\pm 0.24\%$
GS-KAN $\mathbf{[784,15,15,10]}$ 12,346 $\mathbf{87.03\pm 0.32\%}$
Analysis.
Despite having fewer parameters, GS-KAN outperforms the MLP baseline by a margin of $\mathbf{+1.03\%}$. The consistent performance gap across all seeds indicates that the learnable B-spline activations provide a richer per-parameter expressivity than standard ReLU neurons, even in high-dimensional settings where KANs have traditionally struggled with computational overhead. The results suggest that the layer-wise shared basis strategy effectively scales KAN-like architectures to image vectors while maintaining a memory footprint smaller than comparable MLPs.
V Conclusion and Future Work
In this work, we presented GS-KAN, a parameter-efficient adaptation of Kolmogorov-Arnold Networks grounded in Sprecher’s refinement of the superposition theorem. By utilizing layer-wise shared learnable B-splines and linear transformations, GS-KAN effectively decouples network width from parameter complexity of the learned functions in the network. The results demonstrate that the architecture offers significant value for memory-constrained applications, combining efficiency with robust performance.
Our architecture can be interpreted as a relaxation of Sprecher’s functional representation. Theoretically, the learnable linear parameters ($\lambda_{p,q},\epsilon_{q}$) in GS-KAN encompass the specific constants required by Sprecher’s theorem as a subset of the solution space. Unlike earlier numerical implementations of “Sprecher networks”—which necessitated pathological or fractal-like inner functions to strictly adhere to the theorem, our relaxation distributes the informational burden across these generalized linear weights. This structural flexibility suggests that smooth B-spline approximations are sufficient for high-precision learning within this framework, effectively bridging the gap between exact representation theory and practical gradient-based optimization.
Our experiments show that this generalized approach achieves:
  1. 1.
    Superior Function Approximation: GS-KAN outperforms MLPs by orders of magnitude ($10^{-4}$ vs $10^{-2}$ MSE) on continuous functions and demonstrates higher expressivity per parameter than the other tested KAN architectures.
  2. 2.
    Scalability in High Dimensions: In image classification tasks ($D_{in}=784$), GS-KAN surpasses MLPs (+1.03% accuracy) and enables spline-based learning in strict parameter regimes where Standard KANs are computationally infeasible due to grid scaling.
  3. 3.
    Competitive Real-World Performance: On tabular benchmarks, GS-KAN matches the performance of current KAN variants, often yielding superior accuracy-per-parameter ratios in low-resource settings.
Future Work.
While GS-KAN demonstrates promising results, its full potential as a novel architecture remains to be explored. Future research directions include:
  • Adaptive Grid Distributions: Currently, GS-KAN uses fixed uniform spline grids. Implementing learnable knot positions could allow the model to concentrate resolution on complex regions of the function, significantly boosting efficiency against fixed-grid baselines.
  • Computationally Efficient Basis Functions: While B-splines offer excellent expressivity through local plasticity, their recursive evaluation can be computationally intensive during training. Future work could explore faster alternatives with similar local properties, such as Gaussian Radial Basis Functions (RBFs), to reduce training time and memory overhead without sacrificing the model’s adaptability.
  • Specialized Optimization: Our benchmarks utilized standard optimizers (Adam) and heuristic initialization. Developing initialization schemes and learning rate schedules specifically tailored to the coupled dynamics of linear weights and spline coefficients could further widen the performance gap against established baselines.
  • Deep Residual Architectures: Extending GS-KAN with residual connections could enable the training of significantly deeper networks, potentially unlocking performance gains on large-scale datasets.